EACH MULTIPLE CHOICE QUESTION IS WORTH 4 POINTS 1) Four forces of the same magnitude act on a rectangle that can rotate about explain your reasoning
Answer #1

We know torque \vec{T}=\vec{r}\times \vec{F}

  or\: \: \: \: \: \vec{T}=\left | \vec{r} \right |.\left | \vec{F} \right |\sin \Theta

For this problem \Theta =90o.

  \therefore \: \: \: \: \: \vec{T}=\left | \vec{r} \right |.\left | \vec{F} \right |\: \: \: \: \: \: \: \: (i)

Here r is the perpendicular distance and F is the magnitude of the applied force.

Let the perpendicular distance for F1=r1 .

Let the perpendicular distance for F2=r2.

     Let the perpendicular distance for F3=r3.

   Let the perpendicular distance for F4=r4.

From the figure we can say

r_{4}> r_{2}> r_{1}> r_{3}\: \: \: \: \: \: \left [ \: \: \: r_{3}=0\: \: \right ]

From the equation (i) we can say option  

     (E)\: \: \: \: \: \: \: T_{4}> T_{2}> T_{1}> T_{3}\: \: \: \: \: \: \left [ \: \: \: T_{3}=0\times F_{3}=0\: \: \right ]

is correct.